OpenStudy (anonymous):

f(x)=10-x f-1(x)=____________ top is neg power of f but couldnt write -1 power of f

OpenStudy (anonymous):

Inverse it. y = 10 -x is your original equation Switch the x and y. x = 10 - y Move y over to get the original form. 10 - x = y

OpenStudy (anonymous):

\[f^{-1}(x) = 10-x\]We check this as: \[f^{-1}(f(x)) = 10-(10-x)=x\]

OpenStudy (anonymous):

well if you are finding the inverse you set the function as y=10-x then swap your x and y's x=10-y then solve for y